solve the equation w+(3)=9
Answer by bucky(2189)

Given the equation:
.
w + (3) = 9
.
To solve this equation means to find the value of w that will make both sides of this
equation equal to each other.
.
For this equation you can ask yourself, "What number when I add three to it will equal 9?"
.
In asking this question you may be able to do the math in your head to come up with the
answer that if w was 6 you could add 3 to it and the answer would be 9, the same as the
right side of the equation.
.
But to solve the equation using a more general approach we need to get w by itself on
one side of the equation and everything else on the other side. Then we will know what
w is equal to.
.
So what we need to do in this problem is to get rid of the + 3 on the left side so that
we just have w by itself. We can do this by subtracting 3 from the left side. But in
equations, whatever we do to one side, we also have to do to the other side to keep the
equation in balance. So if we subtract 3 from one side we must also subtract 3 from the
other side. When we subtract 3 from the left side and from the right side, the equation
becomes:
.
w + 3 - 3 = 9 - 3
.
Notice on the left side that the + 3 and the - 3 cancel each other out by becoming
zero. Therefore, on the left side we are just left with w. On the right side the
subtraction of 3 from 9 results in 6. And what we are left with is just:
.
w = 6
.
And that's the answer to the problem. You can check the answer by returning to the original
problem of:
.
w + 3 = 9
.
and replace w with your answer of 6 to get:
.
6 + 3 = 9
.
This obviously is true because 6 + 3 does equal 9. So our answer is correct.
.
One additional comment. Don't let the parentheses in +(3) confuse you. Since they are
preceded by a + sign, you can just erase the parentheses to get + 3. If there were a
-(3) then in removing the parentheses you would get - 3.
.
Hope this helps you to understand the problem.
.